OpenStudy (anonymous):

Rewrite sin 3x with only sin x and cos x

OpenStudy (anonymous):

\[\sin(3x) = \sin(x + 2x)\]

OpenStudy (anonymous):

Nw use the formula and expand...

OpenStudy (anonymous):

\[\sin(x+y) = sinx.cosy + cosx.siny\]

OpenStudy (anonymous):

sin x cos2x - sin3x + cos3x?

OpenStudy (anonymous):

@Yahoo!